OpenStudy (lovelyanna):

Find the constant of variation for the quadratic variation. x 2 3 4 5 6 y 24 54 96 150 216 A. 12 B. 6 C. 30 D. 18

OpenStudy (lovelyanna):

@whpalmer4

OpenStudy (whpalmer4):

Take each value of x and square it (as this is a quadratic variation). Now take each value of y and divide it by the corresponding value of x^2. Do you see a pattern?

OpenStudy (lovelyanna):

Its B? correct?

OpenStudy (whpalmer4):

Let's see: x 2 3 4 5 6 y 24 54 96 150 216 x^2 = 4 9 16 25 36 24/4 = 6 54/9 = 6 96/16 = 6 yep! :-)

OpenStudy (lovelyanna):

I get it now! thanks =)

OpenStudy (lovelyanna):

Write a quadratic variation equation if g(x) varies directly with x2, and g(x) = 108 when x = 6. A. g(x) = 1/3x^2 B. g(x) = 3x^2 C. g(x) = 6x^2 D. g(x) = 16x^2 @whpalmer4

OpenStudy (lovelyanna):

B... :/

OpenStudy (lovelyanna):

Is my answer wrong..?

OpenStudy (whpalmer4):

if g(x) varies directly with \(x^2\) then \(g(x) = kx^2\) where \(k\) is the constant of variation \[g(x) = kx^2\]\[g(6) = 108 = k(6)^2\]\[108=k*36\]\[k=3\]so your answer is correct!

OpenStudy (whpalmer4):

varies directly means constant * thing varying varies indirectly means constant / thing varying varies jointly means constant * things varying

OpenStudy (lovelyanna):

Yay! thanks! can you help me with just one more? graphs.. I'm not soo good with them. I think I know the answer tho..

OpenStudy (whpalmer4):

okay, post a different problem and nudge me via "@whpalmer4" please

OpenStudy (lovelyanna):

ok